OpenStudy (alibaby):

There are (5^3)2 ⋅ 5^0 hens in a bird enclosure. What is the total number of hens in the enclosure? 0 5^5 5^6 5^30

OpenStudy (alibaby):

@johnweldon1993 @rishavraj @Joshiscool

OpenStudy (misty1212):

HI!!

rishavraj (rishavraj):

is it \[(5^3)^2 \times5^0\]

OpenStudy (misty1212):

they really tried hard to make this a "word problem" didn't they?

rishavraj (rishavraj):

by the way u know \[a^0 = 1 ~~so~~5^0 = 1\]

OpenStudy (alibaby):

ya @rishavraj it is that

rishavraj (rishavraj):

and also do u know \[(a^x)^y = a^{xy}\]

OpenStudy (alibaby):

yes XD i think the answer is B or C

OpenStudy (alibaby):

@rishavraj

rishavraj (rishavraj):

why B???